d r? + ind the area of the region shared by Consider the circles r? both circles.
Answer #1

x2 +y2=32 40 x*(4-3)= 3x +(9=)= V4 -( AX 80 Gsx) from 0, x=-% C enyeza put ® in ® . -ye+3+ y - 3y + % = 9/6 1 ) 34 = ² => y =.. Y, (x) = 327594-22 Here the internal [a, b] = [ - ]. .. Equation ③ becomes Shaded Area = J (32-x² (32 + 1 % -22) dx rah m= [ + s) = [] += poly) | zz + y s (JR)-ACS -至-35/7)